### v1

@@ 4.3 Mirror Descent Update
Thus, Algorithm 3 outputs $\mathbf{w}_{e_{1}}^{(t+1)}=\min(\frac{W_{1}\cdot\mathbf{y}^{(t+1)}_{e_{1}}}{S_{1}},\mathbf{u}_{e_{1}})=\mathbf{u}_{e_{1}}$.
@@
Thus, Algorithm 3 outputs $\mathbf{w}_{e_{i}}^{(t+1)}=\min(\frac{W_{i}\cdot\mathbf{y}^{(t+1)}_{e_{i}}}{S_{i}},\mathbf{u}_{e_{i}})=\mathbf{u}_{e_{i}}$, for each $i\leq k$.
@@
Thus, Algorithm 3 outputs $\mathbf{w}_{e_{k+1}}^{(t+1)}=\min(\frac{W_{k+1}\cdot\mathbf{y}^{(t+1)}_{e_{k+1}}}{S_{k+1}},\mathbf{u}_{e_{k+1}})=\exp(-\mu)\cdot\mathbf{y}^{(t+1)}_{e_{k+1}}$, as the last quantity is less than $\mathbf{u}_{e_{k+1}}$.
Now consider any $j\geq k+1$. By induction, it holds that $W_{j}=\sum_{\ell=j}^{N}\mathbf{y}_{e_{\ell}}^{(t+1)}\exp(-\mu)$, $S_{j}=\sum_{\ell=j}^{N}\mathbf{y}_{e_{\ell}}^{(t+1)}$, and thus $W_{j}=\exp(-\mu)S_{j}$. This further implies that Algorithm 3 outputs $\mathbf{w}_{e_{j}}^{(t+1)}=\min(\frac{W_{j}\cdot\mathbf{y}^{(t+1)}_{e_{j}}}{S_{j}},\mathbf{u}_{e_{j}})=\exp(-\mu)\cdot\mathbf{y}^{(t+1)}_{e_{j}}$, as the last quantity is less than $\mathbf{u}_{e_{j}}$.
Therefore, Algorithm 3 always output $\mathbf{w}^{(t+1)}$ such that $\mathbf{w}^{(t+1)}=\mathbf{w}^{*}$. That is, $D_{\Phi}(\mathbf{w}^{(t+1)},\mathbf{y}^{(t+1)})=\min_{\mathbf{w}\in\mathcal{X}}D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$. This finishes the proof of the theorem. ∎
@@ 4.4 Privacy Analysis
The privacy analysis follows from a similar approach to EKKL (20). We will first bound the privacy loss in each inner iteration when computing the noisy gradient $\mathbf{g}^{(t)}$, and apply the advanced composition (Lemma 7) over all steps from Line 3 to Line 11 in Algorithm 2 . Then, we consider the $L$ composition in the outer iteration, hence giving the total privacy guarantee for Algorithm 2.
@@
Recall that $\zeta\sim N(\mathbf{0},\mathbf{I}_{2n})$ is defined in Algorithm 2. Therefore, $(\mathbf{X}^{(t)})^{\frac{1}{2}}\zeta\sim N(\mathbf{0},\mathbf{X}^{(t)})$, $(\mathbf{\overline{X}}^{(t)})^{\frac{1}{2}}\zeta\sim N(\mathbf{0},\mathbf{\overline{X}}^{(t)})$. We will instantiate Lemma 31 with $\mathbf{X}^{*}=\overline{\mathbf{X}}^{(t)}$, $\mathbf{X^{\prime}}^{*}=\mathbf{\overline{X}^{\prime}}^{(t)}$. Now, let us analyze the privacy guarantee of the algorithm.
@@
Algorithm 2 with parameter $\lambda=\Theta(\varepsilon^{-1})\tilde{\ell}_{3}(\hat{G})\sqrt{T}\log^{\frac{3}{2}}(\frac{T}{\delta})$, $\varepsilon_{0}=O(\frac{\varepsilon}{\sqrt{T\log\frac{T}{\delta}}})$ is $(\varepsilon,\delta)$-differentially private.
@@
We choose $\overline{\delta}=\frac{\delta}{2T}$. Then by advanced composition (Lemma 8), the privacy guarantee of $T$ mirror descent steps from Line 3 to Line 11 in Algorithm 2 is $\left(\overline{\varepsilon}\cdot 4\sqrt{T\log\frac{4}{\delta}},\delta\right)$. Additionally, by Line 6 in Algorithm 2, we $(\varepsilon_{4},0)$-differentially privately release $\nu$ each outer iteration $\ell$. To sum up, by composition (Lemma 7), the total privacy guarantee in the update process is $\left(L\overline{\varepsilon}\cdot 4\sqrt{T\log\frac{4}{\delta}}+L\varepsilon_{4},\delta\right)$.
Recall that the release of $W$, $\mathbf{u}$, and ~$\tilde{\ell}_{3}(\hat{G})$ is $(\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3})$-differentially private according to Appendix B. Since $L=\log_{3}(\frac{3}{\beta})$, $\varepsilon_{1}=\varepsilon_{2}=\varepsilon_{3}=\frac{\varepsilon}{6}$ and $\varepsilon_{4}=\frac{\varepsilon}{6L}$, by adaptive composition (Lemma 7), Algorithm 2 is $\left(2\varepsilon/3+L\overline{\varepsilon}\cdot 4\sqrt{T\log\frac{4}{\delta}},\delta\right)$-differentially private, where $\overline{\varepsilon}=O(\frac{1}{\lambda}\ell_{3}(\overline{G})\log\frac{2T}{\delta})$.
Additionally, recall that $\ell_{3}(\overline{G})=\Theta(\ell_{3}(\hat{G}))$ and ~$\tilde{\ell}_{3}(\hat{G})$ is differentially privately released from $\ell_{3}(\hat{G})$. Therefore, when we set $\lambda=\Theta(\varepsilon^{-1})\tilde{\ell}_{3}(\hat{G})\sqrt{T}\log^{\frac{3}{2}}(\frac{T}{\delta})\log(\frac{3}{\beta})$, we can guarantee that the algorithm is $(\varepsilon,\delta)$-DP. ∎
@@
Now we consider the standard notion of edge privacy (Definition 5), that two graphs are called neighboring if the two vectors corresponding to their edge weights differ by at most $1$ in the $\ell_{1}$ norm. We first consider the graph $\overline{G}$ and its neighboring graph. Let $e_{1},e_{2},\dots,e_{\binom{n}{2}}$ denote an ordering of all possible edges. Denote $\overline{G}^{\prime}$ (with edge weight function $\overline{\mathbf{w}}^{\prime}$) as a neighboring graph of $\overline{G}$ (with edge weight function $\overline{\mathbf{w}}$). Then we denote $G_{0}=\overline{G}$, and for $i=1,\dots,\binom{n}{2}$, we let $G_{i}$ denote the graph generated by replacing the weight of edge $e_{i}$ in $G_{i-1}$ with $\overline{\mathbf{w}}_{e_{i}}^{\prime}$. It follows that $G_{\binom{n}{2}}=\overline{G}^{\prime}$. Note that $G_{i-1}$ and $G_{i}$ differ only by $\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|$ at $e_{i}$ entry, and $\sum_{i=1}^{\binom{n}{2}}\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|=1$. We let $\mathbf{A}_{\triangle,i}$ denote the adjacency matrix of triangle-motif graph of $G_{i}$. We have that $\left\|\mathbf{A}_{\triangle,i}-\mathbf{A}_{\triangle,i-1}\right\|_{1}=\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|\cdot\left\|\mathbf{D}_{\triangle,i}^{(e_{i})}\right\|_{1}\leq\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|\cdot\ell_{3}(G_{i})$. In the non-degenerate case (see Corollary 42), we have that $\ell_{3}(\overline{G})=\Theta(\ell_{3}(\hat{G}))=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$; furthermore, we have $\left|\ell_{3}(G_{i})-\ell_{3}(\overline{G})\right|\leq w_{\max}$, which implies that $\ell_{3}(G_{i})=\Theta(\ell_{3}(\overline{G}))$. Thus $\left\|\overline{\mathbf{A}}_{\triangle}-\overline{\mathbf{A}}_{\triangle}^{\prime}\right\|_{1}\leq\sum_{i=1}^{\binom{n}{2}}\left\|{\mathbf{A}}_{\triangle,i}-{\mathbf{A}}_{\triangle,i-1}\right\|_{1}\leq O(\sum_{i=1}^{\binom{n}{2}}\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|\cdot\ell_{3}(\hat{G}))\leq O(\ell_{3}(\hat{G}))$. The remaining proof follows similarly to the analysis in Theorem 32.
@@
Now let us recall that the real input graph is $\hat{G}$ instead of $\overline{G}$. While the distance between the actual input graph $\hat{G}$ and its neighboring graph is at most $1$ (in terms of edge differences), the distance between the corresponding rescaled neighboring graph and $\overline{G}$ is not necessarily $1$. Since in the non-degenerate case it holds that $W=\Omega(\frac{\log(1/\beta)}{\varepsilon})$ and $\overline{G}=\frac{\overline{W}}{\hat{W}}\hat{G}$, it follows that $\overline{G}$ is a scaled version of $\hat{G}$ by a constant factor. Thus the privacy loss in each inner iteration of Algorithm 2 is still $O(\frac{\ell_{3}(\overline{G})}{\lambda})$.
@@ 4.5 Utility Analysis
Denote $Y=\mathrm{Lap}(3/\varepsilon)+3\log(3/\beta)/\varepsilon$. Then by Lemma 11, we have $0<Y<6\log(3/\beta)/\varepsilon$ with probability at least $(1-\beta/3)$. Note that $\hat{W}=\Omega(\frac{\log(1/\beta)}{\varepsilon})=\Omega(Y)$. Thus we have $W=\Theta(\hat{W})$ and $\frac{Y}{W}=O(1)$. Therefore,
@@
Let $\mathbf{g}^{(t)}$ be the estimated gradient in Algorithm 2. Then we have, $\mathbb{E}_{{\{\zeta,\nu_{e}^{(\ell)}\}}}[\mathbf{g}^{(t)}]=\nabla f_{\triangle}(\mathbf{w}^{(t)})$ and $\mathbb{E}_{{\{\zeta,\nu_{e}^{(\ell)}\}}}[\left\|\mathbf{g}^{(t)}\right\|_{\infty}^{2}]=O\left(\left({U}_{\triangle}+{U}_{\Lambda}L/\varepsilon\right)^{2}\log^{2}n\right)$.
@@
Since $\zeta\sim N(\mathbf{0},\mathbf{I}_{2n})$, for any $i$, $\zeta_{i}\zeta_{i}$ follows the chi–squared distribution with expectation equal to $1$, and for any $i\neq j$, $\zeta_{i}\zeta_{j}$ follows the product normal distribution with expectation equal to 00, i.e., $\mathbb{E}[\zeta\zeta^{\top}]=\mathbf{I}_{2n}$. Moreover, recall that $\tilde{\mathbf{w}}_{e}^{(\ell)}=\overline{\mathbf{w}}_{e}+\nu_{e}^{(\ell)}$, and $\nu_{e}^{(\ell)}$ is distributed from $\mathrm{Lap}(1/\varepsilon_{4})$ which has expectation 0 . Denote $e=(i,j)$. Therefore,
@@
Since $\|\mathbf{D}_{\triangle}^{(i,j)(t)}\|_{1}\leq 3\sum_{\ell\neq i,j}\mathbf{w}^{(t)}_{(i,\ell)}\mathbf{w}^{(t)}_{(\ell,j)}\leq 3\sum_{\ell\neq i,j}\mathbf{u}_{(i,\ell)}\mathbf{u}_{(\ell,j)}\leq 3{U}_{\triangle}$. The last second inequality is derived from the assumption in Appendix B. Then, by denoting that $(\begin{array}[]{cc}\mathbf{0}&\mathbf{D}_{\triangle}^{(e)(t)}\\ \mathbf{D}_{\triangle}^{(e)(t)}&\mathbf{0}\end{array})=\sum_{i=1}^{(2n)^{2}}c_{i}\mathbf{E}_{i}$ where $\mathbf{E}_{i}$ has only one single non-zero-entry equal to $1$, we can have that
@@
Therefore, similar to the proof of Theorem 32, we have
@@
Applying Theorem 35, we have
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Now that we have bounded all the parameters, we can bound the additive error of Algorithm 2 in the following. First, we give the bound of $f_{\triangle}(\mathbf{w})$ by utilizing Corollary 20.
@@
Let $\mathbf{w}$ be the output of Algorithm 2 with parameter $\beta^{\prime}=\beta/3$, $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$, where $R=O(\sqrt{W\log n}),B=\Theta\left(\left({U}_{\triangle}+{U}_{\Lambda}L/\varepsilon\right)\log n\right)$. Then with probability at least $31-\beta/$ 3,
@@
Through the above discussion, we have $\rho=\frac{1}{W}$. Now we set $\beta^{\prime}=\beta/3$, $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$ in Algorithm 2. By Theorem 29 and Corollary 20, with probability at least $31-\beta/$ 3, it holds that
@@
This further implies that with probability at least $31-\beta/$ 3,
@@
Let $G$ be the output graph of Algorithm 2 with $T=\Theta(\frac{W({\varepsilon}{U}_{\triangle}+{U}_{\Lambda})}{n\log(n/\delta)\tilde{\ell}_{3}(\hat{G})})$, $\beta^{\prime}=\beta/3$, $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$, where $R=O(\sqrt{W\log n}),B=\Theta\left(\left({U}_{\triangle}+{U}_{\Lambda}L/\varepsilon\right)\log n\right)$. Assume that $\hat{W}=\Omega(\frac{\log(1/\beta)}{\varepsilon})$ and $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$. Then the triangle-motif cut difference between $G$ and the original graph $\hat{G}$ is at most
@@
$h(\mathbf{w})=\max_{\mathbf{X}\in\mathcal{D}^{\prime}}H(\mathbf{w},\mathbf{X}),\ \ \text{ where }H(\mathbf{w},\mathbf{X})=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{A}_{\triangle}-\overline{\mathbf{A}}_{\triangle}\\ \mathbf{A}_{\triangle}-\overline{\mathbf{A}}_{\triangle}&\mathbf{0}\end{array}\right)\bullet\mathbf{X}$
@@
Let $\mathbf{X}^{(2)}$ be the maximizer of $h(\mathbf{w})\in\mathcal{D}^{\prime}$ Let $\mathbf{X}^{(3)}=(1-\frac{1}{n})\cdot\mathbf{X}^{(2)}+\frac{1}{n}\cdot\mathbf{I}_{2n}$. We can see that $\mathbf{X}^{(3)}\in\mathcal{D}$. Then we have,
@@
Recall that, according to Theorem 32, $\lambda=\Theta(\varepsilon^{-1})\tilde{\ell}_{3}(\hat{G})\sqrt{T}\log^{\frac{3}{2}}(\frac{T}{\delta})\log(3/\beta)$ and $L=\log_{3}(\frac{3}{\beta})$. By the assumption $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$, we have ~$\tilde{\ell}_{3}(\hat{G})=\Theta(\ell_{3}(\hat{G}))$. Moreover, note that, the algorithm fails with at most $32\beta/$ 3 probability referring to Section 3. Then we can conclude that, by choosing $T=\Theta(\frac{W({\varepsilon}{U}_{\triangle}+{U}_{\Lambda})}{n\log(n/\delta)\tilde{\ell}_{3}(\hat{G})})$, the triangle-motif cut difference between the output graph $G$ (with edge weights $\mathbf{w}$) and $\hat{G}$ is at most
@@
Let $G$ be the output graph of Algorithm 2 with $T=\Theta(\frac{W({\varepsilon}{U}_{\triangle}+{U}_{\Lambda})}{n\log(n/\delta)\tilde{\ell}_{3}(\hat{G})})$, $\beta^{\prime}=\beta/3$, $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$, where $R=O(\sqrt{W\log n}),B=\Theta\left(\left({U}_{\triangle}+{U}_{\Lambda}L/\varepsilon\right)\log n\right)$. Denote $\hat{W}$ and $w_{\max}$ as the sum of edge weights and the maximum edge weight of the original graph $\hat{G}$ separately. Then the triangle-motif cut difference between $G$ and $\hat{G}$ is at most
@@
We first assume that $\hat{W}=\Omega(\frac{\log(1/\beta)}{\varepsilon})$ and $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$, which we refer to as the non-degenerate case. By Lemma 41, ${U}_{\triangle}=O(nw^{2}_{\max}\frac{\log^{2}(n/\beta)}{\varepsilon^{2}})$, ${U}_{\Lambda}=O(nw_{\max}\frac{\log(n/\beta)}{\varepsilon})$. Therefore, by Theorem 40, the additive error is at most
@@
Finally, we note that the correctness of Theorem 3 follows from Theorem 32 and Corollary 42 with error probability $\beta=\frac{1}{4}$.

### v2

@@ 4.3 Mirror Descent Update
Thus, Algorithm˜3 outputs $\mathbf{w}_{e_{1}}^{(t+1)}=\min(\frac{W_{1}\cdot\mathbf{y}^{(t+1)}_{e_{1}}}{S_{1}},\mathbf{u}_{e_{1}})=\mathbf{u}_{e_{1}}$.
@@
Thus, Algorithm˜3 outputs $\mathbf{w}_{e_{i}}^{(t+1)}=\min(\frac{W_{i}\cdot\mathbf{y}^{(t+1)}_{e_{i}}}{S_{i}},\mathbf{u}_{e_{i}})=\mathbf{u}_{e_{i}}$, for each $i\leq k$.
@@
Thus, Algorithm˜3 outputs $\mathbf{w}_{e_{k+1}}^{(t+1)}=\min(\frac{W_{k+1}\cdot\mathbf{y}^{(t+1)}_{e_{k+1}}}{S_{k+1}},\mathbf{u}_{e_{k+1}})=\exp(-\mu)\cdot\mathbf{y}^{(t+1)}_{e_{k+1}}$, as the last quantity is less than $\mathbf{u}_{e_{k+1}}$.
Now consider any $j\geq k+1$. By induction, it holds that $W_{j}=\sum_{\ell=j}^{N}\mathbf{y}_{e_{\ell}}^{(t+1)}\exp(-\mu)$, $S_{j}=\sum_{\ell=j}^{N}\mathbf{y}_{e_{\ell}}^{(t+1)}$, and thus $W_{j}=\exp(-\mu)S_{j}$. This further implies that Algorithm˜3 outputs $\mathbf{w}_{e_{j}}^{(t+1)}=\min(\frac{W_{j}\cdot\mathbf{y}^{(t+1)}_{e_{j}}}{S_{j}},\mathbf{u}_{e_{j}})=\exp(-\mu)\cdot\mathbf{y}^{(t+1)}_{e_{j}}$, as the last quantity is less than $\mathbf{u}_{e_{j}}$.
Therefore, Algorithm˜3 always output $\mathbf{w}^{(t+1)}$ such that $\mathbf{w}^{(t+1)}=\mathbf{w}^{*}$. That is, $D_{\Phi}(\mathbf{w}^{(t+1)},\mathbf{y}^{(t+1)})=\min_{\mathbf{w}\in\mathcal{X}}D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$. This finishes the proof of the theorem. ∎
@@ 4.4 Privacy Analysis
The privacy analysis follows from a similar approach to EKKL (20). We will first bound the privacy loss in each inner iteration when computing the noisy gradient $\mathbf{g}^{(t)}$, and apply the advanced composition (Lemma 7) over all steps from Line 3 to Line 11 in Algorithm˜2 . Then, we consider the $L$ composition in the outer iteration, hence giving the total privacy guarantee for Algorithm˜2.
@@
Recall that $\zeta\sim N(\mathbf{0},\mathbf{I}_{2n})$ is defined in Algorithm˜2. Therefore, $(\mathbf{X}^{(t)})^{\frac{1}{2}}\zeta\sim N(\mathbf{0},\mathbf{X}^{(t)})$, $(\mathbf{\overline{X}}^{(t)})^{\frac{1}{2}}\zeta\sim N(\mathbf{0},\mathbf{\overline{X}}^{(t)})$. We will instantiate Lemma 31 with $\mathbf{X}^{*}=\overline{\mathbf{X}}^{(t)}$, $\mathbf{X^{\prime}}^{*}=\mathbf{\overline{X}^{\prime}}^{(t)}$. Now, let us analyze the privacy guarantee of the algorithm.
@@
Algorithm˜2 with parameter $\lambda=\Theta(\varepsilon^{-1})\tilde{\ell}_{3}(\hat{G})\sqrt{T}\log^{\frac{3}{2}}(\frac{T}{\delta})$, $\varepsilon_{0}=O(\frac{\varepsilon}{\sqrt{T\log\frac{T}{\delta}}})$ is $(\varepsilon,\delta)$-differentially private.
@@
We choose $\overline{\delta}=\frac{\delta}{2T}$. Then by advanced composition (Lemma 8), the privacy guarantee of $T$ mirror descent steps from Line 3 to Line 11 in Algorithm˜2 is $\left(\overline{\varepsilon}\cdot 4\sqrt{T\log\frac{4}{\delta}},\delta\right)$. Additionally, by Line 6 in Algorithm˜2, we $(\varepsilon_{4},0)$-differentially privately release $\nu$ each outer iteration $\ell$. To sum up, by composition (Lemma 7), the total privacy guarantee in the update process is $\left(L\overline{\varepsilon}\cdot 4\sqrt{T\log\frac{4}{\delta}}+L\varepsilon_{4},\delta\right)$.
[…]
@@
Now we consider the standard notion of edge privacy (Definition 5), that two graphs are called neighboring if the two vectors corresponding to their edge weights differ by at most 11 in the $\ell_{1}$ norm. We first consider the graph $\overline{G}$ and its neighboring graph. Let $e_{1},e_{2},\dots,e_{\binom{n}{2}}$ denote an ordering of all possible edges. Denote $\overline{G}^{\prime}$ (with edge weight function $\overline{\mathbf{w}}^{\prime}$) as a neighboring graph of $\overline{G}$ (with edge weight function $\overline{\mathbf{w}}$). Then we denote $G_{0}=\overline{G}$, and for $i=1,\dots,\binom{n}{2}$, we let $G_{i}$ denote the graph generated by replacing the weight of edge $e_{i}$ in $G_{i-1}$ with $\overline{\mathbf{w}}_{e_{i}}^{\prime}$. It follows that $G_{\binom{n}{2}}=\overline{G}^{\prime}$. Note that $G_{i-1}$ and $G_{i}$ differ only by $\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|$ at $e_{i}$ entry, and $\sum_{i=1}^{\binom{n}{2}}\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|=1$. We let $\mathbf{A}_{\triangle,i}$ denote the adjacency matrix of triangle-motif graph of $G_{i}$. We have that $\left\|\mathbf{A}_{\triangle,i}-\mathbf{A}_{\triangle,i-1}\right\|_{1}=\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|\cdot\left\|\mathbf{D}_{\triangle,i}^{(e_{i})}\right\|_{1}\leq\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|\cdot\ell_{3}(G_{i})$. In the non-degenerate case (see Corollary˜42), we have that $\ell_{3}(\overline{G})=\Theta(\ell_{3}(\hat{G}))=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$; furthermore, we have $\left|\ell_{3}(G_{i})-\ell_{3}(\overline{G})\right|\leq w_{\max}$, which implies that $\ell_{3}(G_{i})=\Theta(\ell_{3}(\overline{G}))$. Thus $)\left\|\overline{\mathbf{A}}_{\triangle}-\overline{\mathbf{A}}_{\triangle}^{\prime}\right\|_{1}\leq\sum_{i=1}^{\binom{n}{2}}\left\|{\mathbf{A}}_{\triangle,i}-{\mathbf{A}}_{\triangle,i-1}\right\|_{1}\leq O(\sum_{i=1}^{\binom{n}{2}}\left|\overline{\mathbf{w}}^{\prime}_{e_{i}}-\overline{\mathbf{w}}_{e_{i}}\right|\cdot\ell_{3}(\hat{G}))\leq O(\ell_{3}(\hat{G}$)). The remaining proof follows similarly to the analysis in Theorem˜32.
@@
Now let us recall that the real input graph is $\hat{G}$ instead of $\overline{G}$. While the distance between the actual input graph $\hat{G}$ and its neighboring graph is at most $1$ (in terms of edge differences), the distance between the corresponding rescaled neighboring graph and $\overline{G}$ is not necessarily $1$. Since in the non-degenerate case it holds that $W=\Omega(\frac{\log(1/\beta)}{\varepsilon})$ and $\overline{G}=\frac{\overline{W}}{\hat{W}}\hat{G}$, it follows that $\overline{G}$ is a scaled version of $\hat{G}$ by a constant factor. Thus the privacy loss in each inner iteration of Algorithm˜2 is still $O(\frac{\ell_{3}(\overline{G})}{\lambda})$.
@@ 4.5 Utility Analysis
[…]
@@
Let $\mathbf{g}^{(t)}$ be the estimated gradient in Algorithm˜2. Then we have, $\mathbb{E}_{{\{\zeta,\nu_{e}^{(\ell)}\}}}[\mathbf{g}^{(t)}]=\nabla f_{\triangle}(\mathbf{w}^{(t)})$ and $\mathbb{E}_{{\{\zeta,\nu_{e}^{(\ell)}\}}}[\left\|\mathbf{g}^{(t)}\right\|_{\infty}^{2}]=O\left(\left({U}_{\triangle}+{U}_{\Lambda}L/\varepsilon\right)^{2}\log^{2}n\right)$.
@@
Since $\zeta\sim N(\mathbf{0},\mathbf{I}_{2n})$, for any $i$, $\zeta_{i}\zeta_{i}$ follows the chi–squared distribution with expectation equal to $1$, and for any $i\neq j$, $\zeta_{i}\zeta_{j}$ follows the product normal distribution with expectation equal to 0, i.e., $\mathbb{E}[\zeta\zeta^{\top}]=\mathbf{I}_{2n}$. Moreover, recall that $\tilde{\mathbf{w}}_{e}^{(\ell)}=\overline{\mathbf{w}}_{e}+\nu_{e}^{(\ell)}$, and $\nu_{e}^{(\ell)}$ is distributed from $\mathrm{Lap}(1/\varepsilon_{4})$ which has expectation 0 . Denote $e=(i,j)$. Therefore,
@@
Since $\|\mathbf{D}_{\triangle}^{(i,j)(t)}\|_{1}\leq 3\sum_{\ell\neq i,j}\mathbf{w}^{(t)}_{(i,\ell)}\mathbf{w}^{(t)}_{(\ell,j)}\leq 3\sum_{\ell\neq i,j}\mathbf{u}_{(i,\ell)}\mathbf{u}_{(\ell,j)}\leq 3{U}_{\triangle}$. The last second inequality is derived from the assumption in Appendix˜B. Then, by denoting that $(\begin{array}[]{cc}\mathbf{0}&\mathbf{D}_{\triangle}^{(e)(t)}\\ \mathbf{D}_{\triangle}^{(e)(t)}&\mathbf{0}\end{array})=\sum_{i=1}^{(2n)^{2}}c_{i}\mathbf{E}_{i}$ where $\mathbf{E}_{i}$ has only one single non-zero-entry equal to 11, we can have that
@@
Therefore, similar to the proof of Theorem˜32, we have
@@
Applying Theorem˜35, we have
@@
$\mathbb{E}\left[\left\|\mathbf{g}^{(t)}\right\|_{\infty}^{2}\right]$ $=\int_{s=0}^{\infty}s\cdot\mathsf{pdf}_{\left\|\mathbf{g}^{(t)}\right\|_{\infty}^{2}}ds$
@@
$\leq O(K^{2}\log^{2}n)+O(1)\cdot\int_{s=1}^{\infty}K^{2}(s+O(\log^{2}n))^{2}\cdot n^{2}e^{-s-\log n^{2}}ds$
$\leq O(K^{2}\log^{2}n)+O(K^{2}\log^{2}n)\cdot\int_{s=1}^{\infty}e^{-s}ds+O(K^{2})\int_{s=1}^{\infty}e^{-s}s^{2}ds$
@@
Then we bound the parameter $R$ and $\rho$, which are also defined in Corollary 20. Recall that we choose $\left\|\cdot\right\|$ to be $l_{1}$ norm.
@@
A function $f(\mathbf{x})$ is $\rho$-strongly convex with respect to $\left\|\cdot\right\|$ if and only if
@@
Now that we have bounded all the parameters, we can bound the additive error of Algorithm˜2 in the following. First, we give the bound of $f_{\triangle}(\mathbf{w})$ by utilizing Corollary 20.
@@
Let $\mathbf{w}$ be the output of Algorithm˜2 with parameter $\beta^{\prime}=\beta/3$, $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$, where $R=O(\sqrt{W\log n}),B=\Theta\left(\left({U}_{\triangle}+{U}_{\Lambda}L/\varepsilon\right)\log n\right)$. Then with probability at least $1-\beta/3$,
@@
Through the above discussion, we have $\rho=\frac{1}{W}$. Now we set $\beta^{\prime}=\beta/3$, $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$ in Algorithm˜2. By Theorem˜29 and Corollary 20, with probability at least $1-\beta/3$, it holds that
@@
This further implies that with probability at least $1-\beta/3$,
@@
[…]
@@
[…]
@@
Let $\mathbf{X}^{(2)}$ be the maximizer of $h(\mathbf{w})\in\mathcal{D}^{\prime}$ . Let $\mathbf{X}^{(3)}=(1-\frac{1}{n})\cdot\mathbf{X}^{(2)}+\frac{1}{n}\cdot\mathbf{I}_{2n}$. We can see that $\mathbf{X}^{(3)}\in\mathcal{D}$. Then we have,
@@
[…]
@@
[…]
@@
We first assume that $\hat{W}=\Omega(\frac{\log(1/\beta)}{\varepsilon})$ and $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$, which we refer to as the non-degenerate case. By Lemma 41, ${U}_{\triangle}=O(nw^{2}_{\max}\frac{\log^{2}(n/\beta)}{\varepsilon^{2}})$, ${U}_{\Lambda}=O(nw_{\max}\frac{\log(n/\beta)}{\varepsilon})$. Therefore, by Theorem˜40, the additive error is at most
@@
Finally, we note that the correctness of Theorem˜3 follows from Theorem˜32 and Corollary 42 with error probability $\beta=\frac{1}{4}$.